OpenStudy (anonymous):

what is the height and surface area of a cylinder if the radius is 4cm n volume 350cm3

OpenStudy (mertsj):

\[V=\pi r^2h\]

OpenStudy (mertsj):

Plug into that and find the height. Can you do that?

OpenStudy (anonymous):

h = V/ πr^2 = 350 /( 16 * 3.14 ) = 6.97 cm

OpenStudy (anonymous):

tanx much now can u help me find the surface area

OpenStudy (mertsj):

Well, chlorophyll did the first part for you so she probably wants to do the second part as well.

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

chorophyll can u help me with the surface area plz

OpenStudy (anonymous):

mertj can u help me

OpenStudy (mertsj):

I'm on the phone. When I get off, I will help you and as soon as I start to help you chlorophyll will show up.

OpenStudy (anonymous):

tanx

OpenStudy (anonymous):

people can urll help me now cause the project is due on monday n am just startin the project so can urll help me quickly as possible

OpenStudy (mertsj):

Ok. Think of the surface of the cylinder. It has a circular top and bottom. so find the area of the two circles and tell me what you get.

OpenStudy (anonymous):

how can i find it n i dnt no the formula to find the surface area

OpenStudy (mertsj):

What is the formula for the area of a circle?

OpenStudy (mertsj):